OpenStudy (anonymous):

An open-topped box can be created by cutting congruent squares from each of the four corners of a piece of cardboard that has deminesions of 20cm by 30cm and folding up the sides. Determine the dimennsions of the squares that must be cut to create a box with a volume of 1008cm^3. length=30-2x width=20-2x h=x I dont get this because x and cm arent interchangeable digits

OpenStudy (anonymous):

you would multiply the length, width, and height together and set it equal to 1008cm^3

OpenStudy (anonymous):

(30-2x)(20-2x)(x) = 1008cm^3

OpenStudy (anonymous):

x(30-2x)(20-2x) = 1008

OpenStudy (anonymous):

x(600-60x-40x+4x^2)=1008 x(4x^2-100x+600)=1008 4x^3-100x^2+600x-1008=0 factor out 4; 4(x^3-25x^2+150x-250)=0 divide by 4 on both sides; x^3-25x^2+150x-250=0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome